### v3

# ESC: Efficient Speech Coding with Cross-Scale Residual Vector Quantized Transformers

Abstract
Neural speech codecs aim to compress input signals into minimal bits while maintaining content quality in a low-latency manner. However, existing neural codecs often trade model complexity for reconstruction performance. These codecs primarily use convolutional blocks for feature transformation, which are not inherently suited for capturing the local redundancies in speech signals. To compensate, they require either adversarial discriminators or a large number of model parameters to enhance audio quality. In response to these challenges, we introduce the Efficient Speech Codec (ESC) [^1] Code and pretrained models available at https://github.com/yzGuu830/efficient-speech-codec, a lightweight, parameter-efficient speech codec based on a cross-scale residual vector quantization scheme and transformers. Our model employs mirrored hierarchical window transformer blocks and performs step-wise decoding from coarse-to-fine feature representations. To enhance bitrate efficiency, we propose a novel combination of vector quantization techniques along with a pre-training paradigm. Extensive experiments demonstrate that ESC can achieve high-fidelity speech reconstruction with significantly lower model complexity, making it a promising alternative to existing convolutional audio codecs.
ESC: Efficient Speech Coding with Cross-Scale Residual Vector Quantized Transformers
Yuzhe Gu1,2, Enmao Diao2 1University of Pennsylvania, Philadelphia, PA 2Duke University, Durham, NC tracygu@seas.upenn.eduenmao.diao@duke.edu
1 Introduction
Recent advancements in deep learning have demonstrated the superiority of neural speech codecs over traditional ones, which rely on complex expert design and psycho-acoustic knowledge Valin et al. (2012); Dietz et al. (2015). Early efforts integrating deep generative models, such as WaveNet Oord et al. (2016) and SampleRNN Mehri et al. (2017), into audio codecs have delivered promising results. These models, acting as powerful decoders, synthesize high-quality speech from intermediate representations produced by traditional codecs Kleijn et al. (2018); Klejsa et al. (2019). However, their auto-regressive nature of the decoding process often introduces significant inference latency, limiting their practical application.
Alternatively, some end-to-end neural audio codecs leverage the vector quantization (VQ) network first introduced by Van Den Oord et al. (2017). VQ networks use a learnable collection of code-vectors, known as a codebook, to quantize continuous vectors by assigning them to the nearest codeword. This discretization positions VQNs well-suited for both generation and compression tasks. Following this approach, existing VQ codecs Zeghidour et al. (2021); Défossez et al. (2023); Kumar et al. (2023) typically employ a three-stage architecture: a convolutional encoder and decoder, and a residual vector quantization (RVQ) module Vasuki and Vanathi (2006) applied in the latent space. The encoder and decoder downsample and upsample audio waveform features, creating hierarchical representations. RVQ further refines vanilla vector quantization by minimizing quantization error through a series of codebooks that recursively quantize the residuals from previous stages. Additionally, these codecs employ adversarial discriminators to remove artifacts and produce high-fidelity audio reconstructions. Substantial effort has been dedicated to designing effective audio discriminators, including an improved feature matching loss Kumar et al. (2019), as well as various multi-resolution waveform and spectrogram discriminators Kong et al. (2020); Zeghidour et al. (2021); Défossez et al. (2023); gil Lee et al. (2023). VQ-based audio codecs have demonstrated remarkable performance in audio reconstruction, even at ultra-low bitrates.
Despite these advantages, we find that convolutional VQ codecs heavily depend on powerful discriminators to produce high-quality audio, posing additional optimization challenges due to adversarial training. Moreover, these codecs tend to confront computational constraints, as they require a large number of parameters to balance high compression rates and reconstruction performance. To address these issues, our work develops a more parameter-efficient speech codec by reducing model complexity and implementing the following architectural improvements: 1) replacing convolutional layers with efficient Swin-Transformer Blocks (STBs) Liu et al. (2021), which can better model acoustic features; 2) utilizing the cross-scale residual vector quantization (CS-RVQ) scheme Jiang et al. (2022a) instead of RVQ, extending quantization from a fixed level to multiple levels.
In addition, training VQ codecs frequently leads to a significant challenge: codebook collapse, where a fraction of the codebook remains underutilized in representing input vectors. This issue is frequently observed when training visual tokenizers for generative vision tasks Takida et al. (2022); Zhang et al. (2023); Huh et al. (2023). To address this problem in speech compression, we propose combining product vector quantization (PVQ) Baevski et al. (2019), code factorization Yu et al. (2022), and Euclidean normalization Łańcucki et al. (2020) to enhance codebook utilization. Furthermore, we introduce a learning paradigm to facilitate optimization, which includes a pre-training stage where the codebooks are deactivated and trained subsequently.
In summary, the key contributions of our work are as follows:
  • We introduce ESC, a fully transformer-based speech codec with cross-scale quantization structures. It achieves a superior tradeoff between compression rate, reconstruction quality, and model complexity, outperforming current state-of-the-art models.
  • We propose a novel combination of vector quantization techniques within the cross-scale residual vector quantization (CS-RVQ) framework, coupled with a pre-training paradigm that effectively mitigates codebook collapse and enhances bitrate efficiency.
  • Extensive comparisons with Descript’s audio codec on a multilingual speech corpus demonstrate that transformers and CS-RVQ, the core components of ESC, are superior backbones for speech foundation models than the mainstream convolutions and RVQ.
2 Related Work
2.1 Neural Audio Codecs
Recently, most notable neural audio codecs have been based on the vector quantization (VQ) network, including SoundStream Zeghidour et al. (2021), EnCodec Défossez et al. (2023), and Descript’s audio codec (DAC) Kumar et al. (2023). SoundStream is distinguished as the first universal codec capable of handling diverse audio types. EnCodec improves compression rates by integrating a lightweight transformer language model within the discrete latent space and implements a streaming architecture. Building on similar backbones, Kumar et al. (2023) further explore the implications of quantization dropout, a technique for bitrate scalability, and demonstrate the superiority of periodic inductive bias functions over common activation functions for audio signal modeling gil Lee et al. (2023); Ziyin et al. (2020). These models directly process audio waveforms and are classified as time-domain codecs.
In contrast, frequency-domain codecs focus on processing more intuitive audio spectrogram features. Lyra Kleijn et al. (2021), for example, converts audio waveforms into log mel-spectrograms and directly quantizes them into tokens. Due to the non-invertibility of mel-spectrograms, it relies on a vocoder Kalchbrenner et al. (2018) for waveform synthesis. To circumvent the inefficiencies associated with heavy vocoders, some frequency-domain codecs, including TFNet Jiang et al. (2022b) and our ESC, employ the invertible Short-time Fourier Transform (STFT) to convert waveforms into complex spectra. This design enables the reconstructed STFT spectra to be seamlessly inverted back into waveforms without information loss using inverse-STFT. Among recent audio codecs, DAC achieves state-of-the-art compression ratios and reconstruction quality, though its computation bottlenecks are sometimes overlooked.
2.2 Swin Transformers
Vision Transformers (ViTs) Dosovitskiy et al. (2020) have outperformed convolutional neural networks (CNNs) in various image processing tasks, largely due to their superior ability to capture complex patterns. The Swin Transformer Liu et al. (2021), a notable variant, enhances this capability by employing a hierarchical approach with shifted window attention mechanisms, enabling it to scale efficiently to high-resolution signals while maintaining computational efficiency. In the context of image compression, Swin Transformers have demonstrated exceptional performance. Studies by Zhu et al. (2021) and Zou et al. (2022) show that Swin Transformers surpass CNNs in modeling spatial hierarchies and long-range dependencies. The attention mechanism facilitates the accurate preservation of essential details and textures, even at lower bitrates. These capabilities suggest that transformers could also be effective in applications beyond image compression, such as modeling audio spectrograms.
2.3 Vector Quantization
In the Vector Quantized Variational Autoencoder (VQ-VAE) (Van Den Oord et al., 2017), vector quantization (VQ) functions as a trainable layer that deterministically quantizes encoded latent variables by mapping them to their nearest neighbors in an embedding codebook. A VQ layer, denoted as $Q(\cdot\ ;\mathcal{C})$, is parameterized by a collection of continuous vectors $\mathcal{C}=\{\bm{c}_{1},...,\bm{c}_{K}\}$, each referred to as a codeword, with its associated index known as a code. The layer quantizes a vector $\bm{z}_{e}\in\mathbb{R}^{d}$ to $\bm{z}_{q}\in\mathbb{R}^{d}$ by selecting the Euclidean nearest codeword $\bm{c}_{k}$ from the codebook $\mathcal{C}$, i.e.,
$\bm{z}_{q}:=\bm{c}_{k}=\operatorname*{arg\,min}_{\bm{c}_{j}\in\mathcal{C}}||\bm{z}_{e}-\bm{c}_{j}||_{2}^{2}.$ (1)
For convenience, we denote the output of the VQ function as $(\bm{z}_{q},\tilde{z}_{q})$, where $\tilde{z}_{q}:=k$ represents the discrete code corresponding to the nearest codeword. During compression, the encoding process outputs the discrete index $\tilde{z}_{q}$, which is stored with a $\log_{2}K$ bit budget. The decoding process starts by retrieving the continuous vector $\bm{z}_{q}$ from the codebook using the index $\tilde{z}_{q}$. The VQ function $Q(\cdot;\mathcal{C})$ is non-differentiable due to the $\operatorname*{arg\,min}$ operator. Common strategies use a straight-through estimator (STE) Bengio et al. (2013) to bypass this in back-propagation. In other words, the gradient component $\frac{\partial\bm{z}_{q}}{\partial\bm{z}e}$ is estimated by identity. Additionally, auxiliary losses including codebook loss and commitment loss are proposed to pull the codewords and latent features closer:
$\mathcal{L}_{vq}=||\text{sg}(\bm{z}_{e})-\bm{z}_{q}||_{2}^{2}+\beta||\bm{z}_{e}-\text{sg}(\bm{z}_{q})||_{2}^{2}.$ (2)
Here $\text{sg}(\cdot)$ denotes the stop-gradient operator. The first term updates the codebook with an $l_{2}$ error, pushing the codewords towards the input vectors. The second term ensures that $\bm{z}_{e}$ commits to the embedding without growing arbitrarily. The scalar $\beta$ balances the importance of updating the codebook and the encoder.
2.4 Codebook Collapse
Straight-through estimators (STEs) can lead to significant issues, most notably codebook collapse, as detailed by Vuong et al. (2023). In a recent study, Huh et al. (2023) provide a plausible explanation, attributing the collapse to an internal codebook covariate shift during training. Frequent adjustments in encoder representations cause misalignment with the codebook, resulting in only a subset of codewords being updated. Consequently, VQ layers are prone to divergence, often ending up with a significant number of inactive vectors. Various strategies have been proposed in generative modeling context to address this issue, including stochastic quantization Takida et al. (2022); Zhang et al. (2023), self-annealed soft-to-hard quantization Agustsson et al. (2017), re-initializing codewords using K-means centroids every few epochs Łańcucki et al. (2020); Dhariwal et al. (2020), and reformulating with finite scalar quantization Mentzer et al. (2024). In audio compression, Kumar et al. (2023) address codebook collapse by down-projecting codewords Yu et al. (2022) and normalizing them within a Euclidean ball Łańcucki et al. (2020).
Figure 1: The framework of ESC: input speech is transformed to a complex STFT $\mathcal{X}$ and linearly embedded into patches. Encoder STBs iteratively halve the frequency resolution and produce hierarchical feature representations. Mirrored decoder STBs recover the frequency resolution by progressively leveraging coarse-to-fine quantized residual features between encoder and decoder hidden states. The entire network is solely composed of efficient transformer blocks and vector quantization layers. The figure displays a scenario when the deepest $3$ of $n+1$ total bitstreams (solid lines) are transmitted, with others left inactive.
3 Efficient Speech Codec (ESC)
3.1 Overall Architecture
As illustrated in Figure 1, ESC operates on the complex spectrum $\mathcal{X}\in\mathbb{R}^{2\times F\times T}$ derived from the Short-Time Fourier Transform (STFT) of a speech signal. Here, the real and imaginary components of $\mathcal{X}$ are treated as separate channels. Instead of using strided convolutions, ESC comprises a series of mirrored transformer encoder and decoder layers, each performing downsampling or upsampling to create coarse and fine representations, as described in Section 3.3. Starting from the quantized latents at the bottleneck VQ, the decoder progressively reconstructs the original spectrum by leveraging multi-level quantized residuals between the intermediate features of the encoder and decoder. This cross-scale decoding mechanism is further detailed in Section 3.4. Finally, the reconstructed spectrum $\hat{\mathcal{X}}$ is transformed back into a waveform through the inverse-STFT.
3.2 Notations
We first define some notations for clarity. The encoder and decoder are denoted by $F_{\phi}(\cdot)$ and $G_{\psi}(\cdot)$, respectively, each being a composition of individual layer functions $f_{\phi_{1}},\ldots,f_{\phi_{n}}$ and $g_{\psi_{1}},\ldots,g_{\psi_{n}}$. We use $\mathcal{Z}\in\mathbb{R}^{C\times F\times T}$ to denote a spectrum feature and $\bm{z}\in\mathbb{R}^{CF}$ to denote a flattened time frame vector in $\mathcal{Z}$. Specifically, $\mathcal{Z}_{e_{i}}$ refers to the feature after the $i$-th encoder layer, and $\mathcal{Z}_{q_{i}}$ denotes the $i$-th decoder feature.
$\mathcal{Z}_{e_{i}}=f_{\phi_{i}}\circ...\circ f_{\phi_{1}}\circ\mathcal{Z}_{e_{0}}$ (3)
$\mathcal{Z}_{q_{i}}=g_{\psi_{i}}\circ...\circ g_{\psi_{1}}\circ\mathcal{Z}_{q_{0}},$ (4)
Here, $\mathcal{Z}_{e_{0}}$ is the original input feature and $\mathcal{Z}_{q_{0}}$ is the latent representation at the bottleneck.
3.3 Transformer Encoder and Decoder
To effectively capture redundancies within audio signals, we replace convolutional layers with hierarchical Swin Transformer blocks (STBs) and their extended decoding counterparts.
Patchify. The encoder starts with a linear patchify module, where the complex spectrum $\mathcal{X}$ is divided into small patches and linearly up-projected:
[…]
Here, the patch size across the frequency and temporal dimensions is $(\frac{F}{H_{0}},\frac{T}{W_{0}})$. This step reduces the input resolution to alleviate the computational burden on attention computation. At the end of the decoder, a symmetric de-patchify module reshapes the decoded patch feature $\mathcal{Z}_{q_{n}}$ and linearly down-projects it to produce a recovered spectrum $\hat{\mathcal{X}}$.
Swin Transformer blocks. STBs in both the encoder and decoder employ window-based multi-head self-attention (W-MSA), partitioning spectrum features into smaller windows and computing attention in parallel within each window. This approach enables more efficient computation compared to vanilla attention mechanisms. To ensure connections between windows, STBs cascade two interleaved W-MSAs, with the outputs of the first being shifted for the second. This design allows STBs to capture local and global feature dependencies both effectively and efficiently.
Downsampling and upsampling. ESC maintains temporal resolution while scaling frequency resolution to equalize bitrates across different bitstreams. To achieve this, we modify the original patch merging/splitting modules with a single-dimensional pixel unshuffle/shuffle module Shi et al. (2016) along the frequency dimension. During encoder downsampling, an intermediate encoder spectrum feature $\mathcal{Z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first reshaped and then projected by $P_{e_{i}}\in\mathbb{R}^{vC_{i}\times C_{i+1}}$ as follows:
[…]
where $v$ is the down-scaling factor. The upsampling process mirrors this operation in reverse. An intermediate decoder feature $\mathcal{Z}_{q_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first projected by $P_{q_{i}}\in\mathbb{R}^{C_{i}\times vC_{i+1}}$ and then reshaped, resulting in an up-scaled frequency resolution:
[…]
Overall, the transformer encoder and decoder layers are mirrored, creating symmetric and hierarchical representations of the input audio spectrum. With these backbones, ESC is a fully transformer-based codec without any convolutional modules.
3.4 Cross-Scale Residual Vector Quantization
To achieve parameter-efficient modeling of audio signals, ESC employs multi-scale features that capture coarse-to-fine information. It integrates the more intuitive residual-based cross-scale vector quantization (CS-RVQ) framework proposed by Jiang et al. (2022a), eliminating the need for additional networks to merge encoder and decoder features for improved reconstruction quality. As depicted in Algorithm 1, Algorithm 2 and Figure 1, the decoding process is conditioned on the encoded quantized residuals between encoder and decoder features from low-to-high resolution scales. This approach differs from the commonly used residual vector quantization scheme, which operates solely at the lowest scale, relying on the highest-level information while overlooking low-level details.
Encoding. The encoding process begins with the encoder $F_{\phi}(\cdot)$, creating multi-scale encoder features $\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}$. $\bm{z}_{e_{n}}$ is first quantized by the bottleneck quantizer $Q_{0}$ to form the lowest bitstream. This represents the simplest case when the number of transmitted bitstream $s$ is set to 1, and CS-RVQ reduces to a fixed-scale VQ at the bottleneck. For higher bitstreams, the residual between symmetric encoder and decoder at higher resolutions, $\bm{z}_{e_{n-i+1}}-\bm{z}_{q_{i-1}}$, is quantized by $Q_{i}$. The quantized residual $\bm{q}_{i}$ is then added back to $\bm{z}_{q_{i-1}}$ and decoded by the subsequent decoder layer $g_{\phi_{i}}(\cdot)$, producing the next decoder feature $\bm{z}_{q_{i}}$. Recursively, residuals at higher resolutions are progressively quantized, forming the remaining bitstreams (see Algorithm 1, Lines 3-6). This mechanism enables multi-scale learning, allowing the decoder layers to incrementally reduce quantization errors by conditioning on encoder-decoder residual features. When $s>2$, this encoding process requires forward passing $s-2$ additional decoder layers to produce residuals at higher levels. After encoding, the input $\bm{z}_{e_{0}}$ is compressed into multi-level codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},\ldots,\tilde{z}_{q_{s-1}}$.
Algorithm 1 CS-RVQ Encoding
1:A flattened time frame $\bm{z}_{e_{0}}\in\mathbb{R}^{C_{0}H_{0}}$, encoder $F_{\phi}(\cdot)$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$, number of bitstreams $s$
2:$\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}\leftarrow F_{\phi}(\bm{z}_{e_{0}})$triangleright Encoder forward pass
3:$\bm{z}_{q_{0}},\tilde{z}_{q_{0}}\leftarrow Q_{0}(\bm{z}_{e_{n}})$triangleright bottom VQ
4:for $i=1\dots s-2$ do
5:     $\bm{q}_{i},\tilde{z}_{{q}_{i}}\leftarrow Q_{i}(\bm{z}_{e_{n-i+1}}-\bm{z}_{{q}_{i-1}})$
6:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}}+\bm{q}_{i})$
7:end $\triangleright$ Encoding involves $s-2$ decoder layers
8:if $s>1$ then
9:     $\bm{q}_{s-1},\tilde{z}_{{q}_{s-1}}\leftarrow Q_{i}(\bm{z}_{e_{n-s+2}}\mathrel{-}\bm{z}_{{q}_{s-2}})$
10:end if
11:return $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$
Algorithm 2 CS-RVQ Decoding
1:Codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$
2:$\bm{z}_{q_{0}}\xleftarrow[]{Q_{0}}\tilde{z}_{q_{0}}$triangleright Retrieve codewords from bottom VQ
3:for $i=1\dots s-1$ do
4:     $\bm{q}_{i}\xleftarrow[]{Q_{i}}\tilde{z}_{q_{i}}$
5:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}}+\bm{q}_{i})$
6:end $\triangleright$ Decoding refined by quantized residuals
7:for $i=s\dots n$ do
8:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}})$
9:end $\triangleright$ Continue with regular decoding
10:return $\bm{z}_{q_{n}}$
Decoding. The decoding process starts by retrieving the quantized latent at the bottom VQ using code $\tilde{z}_{q_{0}}$, which provides the initial decoder input $\bm{z}_{q_{0}}$. At higher levels, the codes $\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$ are iteratively used to retrieve codewords, producing multi-scale low-to-high quantized residuals $\bm{q}_{1},...,\bm{q}_{s-1}$. In Algorithm 2, Lines 2-5, each quantized residual $\bm{q}_{i}$ is added back to the corresponding decoder feature $\bm{z}_{q_{i-1}}$ to refine the decoding process. Starting from the $s$-th decoder layer, there are no quantized residuals, and the remaining layers perform regular decoding. Finally, the recovered frame vector $\bm{z}_{q_{n}}$ is obtained, benefiting from $s-1$ quantized residual features.
Training. During training, the encoding and decoding processes are concatenated to form a complete forward pass. To enable bitrate scalability, we sample $s\sim\text{Uniform}\{1,\ldots,n\}$ at a rate $p$ within each training mini-batch. $p$ is a hyperparameter that balances the reconstruction quality at different bitrates, as proposed by Kumar et al. (2023).
3.5 Mitigating Codebook Collapse
ESC performs a per-frame vector quantization. Before nearest neighbor searching, each input spectrum frame feature in $\mathcal{Z}$ needs to be flattened, merging the frequency and channel dimensions. This approach can result in large input vector dimensions for VQ, increasing the optimization challenges associated with codebook underutilization.
Vector quantization setups. To optimize the codebooks effectively, we modify the vanilla VQ by combining product vector quantization with code-vector factorization at each bitstream. Specifically, a flattened $d$-dimensional frame vector $\bm{z}_{e_{i}}$ is split into a set of $l$ sub-vectors. Each sub-vector $\bm{z}_{e_{i}}^{(m)}$ is down-projected by $W_{\text{in}}\in\mathbb{R}^{\frac{d}{l}\times u}$, where $u\ll d$, and then quantized using an individual codebook $\mathcal{C}_{m}$. The selected code-vector is then up-projected by $W_{\text{out}}\in\mathbb{R}^{u\times\frac{d}{l}}$ to form ${\bm{z}}_{q_{i}}^{(m)}$:
$\bm{z}_{e_{i}}\equiv\{\bm{z}_{e_{i}}^{(m)}\mid\bm{z}_{e_{i}}^{(m)}\in\mathbb{R}^{\frac{C_{i}H_{i}}{l}},m=1,...,l\},$ (8)
${\bm{z}}_{q_{i}}^{(m)}=W_{\text{out}}^{\top}\operatorname*{arg\,min}_{\bm{c}_{j}\in\mathcal{C}_{m}}||W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}-\bm{c}_{j}||_{2}.$ (9)
Additionally, both the projected vector $W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}$ and codebook $\mathcal{C}_{m}$ are $l_{2}$ normalized before computing the distance matrix. This equalizes the scales of input vectors and codewords, enhancing codebook optimization by allowing a larger subset of codewords to receive gradients Łańcucki et al. (2020).
Pre-training paradigm. Training transformers can be challenging, and jointly training them with VQ layers is even more difficult. To address this, we propose a pre-training paradigm that includes a warm-start to facilitate the learning process. Initially, all VQ layers are deactivated, meaning no quantization occurs. During this "pre-training" stage, only the encoder and decoder are updated within the CS-RVQ framework, allowing latent features to bypass the quantizers and flow directly into the decoder layers. Once the encoder and decoder have converged by minimizing reconstruction objectives, we resume training the entire VQ codec as usual. This approach helps mitigate the distribution shift of encoder representations by pre-optimizing the encoder. It helps stabilize codebook training and improve bitrate efficiency. Moreover, pre-training an auto-encoder is simpler, as it avoids the quantization errors associated with VQs. The detailed algorithm is provided in Appendix A.
3.6 Training Objectives
To train our codec, we use a combination of reconstruction loss $\mathcal{L}_{recon}$ and vector quantization loss $\mathcal{L}_{vq}$. The reconstruction loss, $\mathcal{L}_{recon}$, consists of two components: an $l_{2}$ distance between the complex spectrum $\mathcal{X}$ and its reconstruction $\hat{\mathcal{X}}$, which forces the model to reconstruct the real and imaginary parts, weighted by $\lambda_{1}$, and a multi-scale mel-spectrogram loss (Kumar et al., 2023), weighted by $\lambda_{2}$. These are denoted as $\mathcal{L}_{stft}$ and $\mathcal{L}_{mel}$:
$\mathcal{L}_{recon}=\lambda_{1}\mathcal{L}_{mel}+\lambda_{2}\mathcal{L}_{stft}.$ (10)
$\mathcal{L}_{vq}$ comprises the standard codebook and commitment losses as described in Equation 2. It is averaged across the $l$ product vector quantizers and summed over all $s$ bitstreams. The final objective for joint optimization is the summation of $\mathcal{L}_{recon}$ and $\mathcal{L}_{vq}$. To deactivate the VQ layers during the pre-training stage, $\mathcal{L}_{vq}$ is set to zero.
Figure 2: Reconstruction quality evaluation of different baseline codecs: dashed lines represent DAC baselines and solid lines represent our ESC models, with x-axis being transmission bits per second and y-axis being PESQ ()($\uparrow$)( ↑ ), Mel-Distance ()($\downarrow$)( ↓ ) and SI-SDR()($\uparrow$)( ↑ ). The metrics are averaged over our composed 1158 10-second speech clips.
4 Experiments
4.1 Experimental Setup
Datasets. We extract 150 hours of 16kHz multilingual clean speech from the DNS Challenge dataset Reddy et al. (2021). Training samples are clipped into 3-second segments, and validation samples into 10-second segments. For evaluation, we compile 1158 multilingual 10-second speech clips with non-overlapping speakers from the LibriSpeech Panayotov et al. (2015), Multilingual LibriSpeech Pratap et al. (2020), and AIShell Shi et al. (2020) datasets.
Baselines. We compare our ESC against the current state-of-the-art time-domain codec DAC, by reproducing three versions [^2] Reproduction settings are detailed in Appendix B.1. on our dataset:
1) DAC-Base (adversarial): Descript’s original released codec, operating on 16kHz audio signals. It has 74M parameter count in total. Its associated discriminator has 42M additional parameter count.
2) DAC-Tiny (adversarial): A smaller version of DAC-Base, with reduced encoder and decoder dimensions, for a fair comparison with ESC.
3) DAC-Tiny (non-adversarial): A smaller and non-adversarial version of DAC to assess the impact of discriminators on improving audio fidelity.
Implementation details. Similar to DAC baselines, we provide different versions of ESC [^3] Complete configurations are detailed in Appendix B.2.:
1) ESC-Base (non-adversarial): A base version codec consisting of 6 encoder/decoder layers, with bitrates ranging from 1.5 to 9.0 kbps. It contains 8.39M parameters when operating at 9.0 kbps.
2) ESC-Base (adversarial): An adversarial version using the same multi-scale multi-band waveform and spectrogram discriminator in DAC.
3) ESC-Large (non-adversarial): A scaled-up version with increased Swin Transformer layer depth, having 15.58M parameters at 9.0 kbps.
Our ESC variants are trained using the AdamW optimizer Loshchilov (2017) with a learning rate of 1e-4 and a weight decay of 1e-2. Training runs up to 0.4 million iterations without learning rate schedulers. The proposed pre-training phase consists of 0.75 million iterations. After pre-training, the codebooks are initialized with a Kaiming normalization distribution He et al. (2015). The quantization dropout rate $p$ is set to 0.75. Loss weighting hyperparameters are set as $\lambda_{1}=0.25$, $\lambda_{2}=1.0$, and the commitment loss weighting $\beta=0.25$. For ESC-Base (adversarial), the $\mathcal{L}_{stft}$ component is eliminated. We use the HingeGAN Lim and Ye (2017) adversarial loss formulation and the $l_{1}$ feature matching loss Kumar et al. (2019), following the approach of DAC.
Automatic evaluation metrics. We use objective metrics to efficiently evaluate reconstruction performance. These include the PESQ score Union (2007) from the speech enhancement domain, following Jiang et al. (2022a); the $l_{1}$ distance between log mel-spectrograms of reference and decoded waveforms (Mel-Distance) Kumar et al. (2023); and the scale-invariant source-to-distortion ratio (SI-SDR) Le Roux et al. (2019). To measure codec inference latency, we use the real-time factor (RTF), defined as the ratio of speech audio duration to model processing time Défossez et al. (2023).
Real Time Factor $\uparrow$
Codec Bitrate #Param. Enc. Dec.
ESC-Base 3.0 kbps 8.10M 33.66 34.97
6.0 kbps 8.21M 27.84 33.02
9.0 kbps 8.39M 24.45 33.95
DAC-Tiny 3.0 kbps 7.96M 42.26 49.52
6.0 kbps 8.07M 44.66 48.63
9.0 kbps 8.17M 43.00 49.10
ESC-Large 3.0 kbps 15.30M 17.91 20.81
6.0 kbps 15.41M 15.48 19.87
9.0 kbps 15.58M 13.73 20.56
DAC-Base 3.0 kbps 73.99M 12.77 3.36
6.0 kbps 74.15M 11.43 3.13
9.0 kbps 74.31M 11.81 3.25
Table 1: Complexity evaluation results of different baseline codecs: RTFs are measured from 100 10-second speech clips on an Intel Xeon Platinum 8352V CPU.
4.2 Comparison with DAC
We provide a thorough comparison focusing on compression rate, reconstruction quality, and inference efficiency, as shown in Table 1 and Figure 2.
Performance evaluation. First, it is important to note that ESC-Base and DAC-Tiny are similar in model size, each with approximately 8 million trainable parameters. Our results show that ESC-Base consistently outperforms DAC-Tiny across all bitrates, even without an adversarial discriminator. In contrast, DAC-Tiny’s reconstruction quality significantly drops without a discriminator in training, particularly in SI-SDR statistics. This indicates a heavy reliance of DAC models on GANs for maintaining high reconstruction quality. Notably, ESC-Base is compatible with the same convolution-based GAN discriminator used in DAC, as evidenced by its improved performance across all metric curves in its adversarial variant. Additionally, ESC-Large demonstrates that increasing ESC’s model size can further enhance performance, with its PESQ curve matching that of DAC-Base, the top-performing and largest model. While DAC-Base achieves higher SI-SDR values, ESC-Large records a smaller Mel-Distance. Thus, we conclude that the two codecs achieve comparable performance, even though ESC-Large is trained without an adversarial discriminator.
Complexity evaluation. Despite their exceptional performance, Descript’s top-performing codecs face significant computational challenges. This is evident from Table 1, where the decoding real-time factor (RTF) for DAC-Base is approximately 3.0, making it rather impractical for real-time applications. In contrast, our transformer-based ESC achieves much higher decoding RTFs (approximately 34), indicating superior computational efficiency. Although ESC-Base is not as fast as DAC-Tiny due to the overhead of attention computation, it offers substantially better speech reconstruction capabilities, striking a favorable balance between compression performance and computation latency. Future work could incorporate transformer speedup techniques such as FlashAttention Dao et al. (2022) to further enhance ESC’s latency further. Moreover, following the CS-RVQ scheme, ESC possesses faster encoding speeds at lower bitrates—a capability not evidently found in DAC models.
These results suggest that our transformer-based codec, equipped with CS-RVQ, is a more parameter-efficient foundation model compared to time-domain convolutional counterparts. ESC is shown to be a more lightweight and effective neural speech codec, as ESC-Large achieves comparable performance to DAC-Base without the need for a powerful discriminator. Specifically, it boasts approximately $\times$4.8 smaller model size, $\times$1.4 faster encoding speed, and $\times$6.4 faster decoding speed.
Method Bitrate PESQ $\uparrow$ Mel dist. $\downarrow$ SI-SDR $\uparrow$ VQ util. $\uparrow$
CNN + RVQ 3.0 kbps 2.71 2.82 0.57 96.8%
6.0 kbps 2.93 2.69 1.03 98.2%
9.0 kbps 2.96 2.68 1.05 98.7%
CNN + CS-RVQ 3.0 kbps 2.70 2.81 2.19 96.6%
6.0 kbps 3.47 2.41 3.79 97.7%
9.0 kbps 3.75 2.25 4.16 97.3%
SwinT + RVQ 3.0 kbps 2.97 2.22 0.77 98.1%
6.0 kbps 3.14 2.08 1.35 99.0%
9.0 kbps 3.16 2.07 1.39 99.2%
ESC-Base (SwinT + CS-RVQ) 3.0 kbps 3.07 2.21 3.55 97.8%
6.0 kbps 3.73 1.80 4.74 98.3%
9.0 kbps 3.92 1.62 5.33 97.9%
ESC-Base w/o Pre-training 3.0 kbps 3.09 2.25 1.75 97.7%
6.0 kbps 3.53 1.97 2.87 98.1%
9.0 kbps 3.58 1.89 2.88 86.5%
Table 2: Performance evaluation of different ablation models: results are obtained from the 1157 10-second speech clips in our test dataset.
4.3 Ablation Study
To investigate the effectiveness of the proposed components in ESC, we conducted thorough ablation experiments [^4] Implementation setups are detailed in Appendix B.3. by training frequency-domain codecs operating on complex STFT spectra with different architectures. For fair comparisons, all other ablation models listed in Table 2 have similar model sizes to ESC-Base.
Swin Transformers and CNNs. To demonstrate that transformers are superior auto-encoder backbones in neural speech coding, we focus on two pairs of experiments: CNN/SwinT + RVQ and CNN/SwinT + CS-RVQ. In these experiments, the channel dimensions of the CNN blocks are set to match the hidden dimensions of the Swin Transformer Blocks (STBs). The comparison, as shown in Table 2, reveals that transformer-based codecs consistently outperform CNN-based codecs across all performance metrics and bitrates, regardless of the quantization scheme used.
CS-RVQ and RVQ. Table 2 highlights that CS-RVQ is a superior quantization scheme compared to RVQ, regardless of whether the backbone is CNN or STB. RVQ-based codecs hit performance bottlenecks, as adding more VQs does not improve audio quality (e.g., from 6.0 kbps to 9.0 kbps). However, codecs using the CS-RVQ scheme do not face such bottlenecks at higher bitrates and consistently outperform their RVQ counterparts. CS-RVQ is therefore a superior vector quantization framework that leverages multi-scale features effectively.
Effect of pre-training paradigm. To evaluate the efficacy of the pre-training stage, we conducted an experiment of ESC-Base w/o pre-training. We monitored the VQ utilization rate, calculated as the sum of entropy (in bits) divided by the maximum number of bits from all transmitted bitstreams. This metric reflects bitrate efficiency and the fraction of seldom-used codewords. The results indicate that models with pre-training achieve a near 1.0 utilization rate. However, ESC-Base w/o pre-training displays a lower utilization rate at 9.0 kbps, and its reconstruction performance is also inferior to that of the fully pre-trained ESC-Base. These findings suggest that the pre-training paradigm indeed helps avoid bitrate wastage and improve audio reconstruction quality.
5 Conclusions
In this paper, we introduce ESC, the first fully transformer-based neural speech foundation model designed for multilingual speech coding. ESC surpasses existing state-of-the-art time-domain VQ-based codecs in terms of complexity and achieves comparable compression performance without the need for a powerful adversarial discriminator. Our extensive evaluations demonstrate that the cross-scale residual vector quantization scheme and the Swin Transformer backbones are better suited for neural speech coding than the convolutional blocks and residual vector quantization utilized in mainstream codecs. Overall, our study suggests a promising direction for speech foundation models. Future research could focus on expanding multi-scale vector quantization techniques and investigating additional transformer variants optimized for speech signal modeling.
6 Limitations
First, recent neural audio codecs are increasingly utilized in downstream generation tasks, where the codec acts as a foundation model to create discrete acoustic representations Borsos et al. (2023); Kreuk et al. (2022); Siuzdak (2023); Wang et al. (2023); Du et al. (2024). These compressed representations, treated as acoustic tokens, are suitable for auto-regressive language modeling in generative tasks. However, our work does not explore this important aspect. A promising future direction would be to evaluate ESC in downstream applications such as speech synthesis and speech recognition. We anticipate that the cross-scale code representations learned from transformer backbones could offer advantages over the fixed-scale features of mainstream convolutional codecs in these tasks.
Second, different automatic metrics for audio evaluation can produce inconsistent results, which is evidenced in our results. To further strengthen our conclusions, it is necessary to conduct subjective evaluations involving human evaluators, such as MUSHRA listening tests Series (2014). Despite this limitation, we provide a collection of demo speech samples publicly available in our codebase, which we hope will help demonstrate ESC’s performance and compensate for the absence of subjective metrics.
Besides, the primary focus of this work is to demonstrate the superiority of transformer and cross-scale frameworks over other mainstream methods, rather than to develop a production-ready codec like DAC or EnCodec. Nonetheless, given the scalability of transformers Kaplan et al. (2020), increasing the ESC model size and training it on larger and more diverse audio datasets also represent a promising direction for enhancing its practical applicability.
Finally, as discussed in Section 3.4, the cross-scale residual vector quantization (CS-RVQ) scheme requires the partial use of decoder layers during the encoding process, introducing additional latency as the bitrate increases. Similar to residual vector quantization, CS-RVQ requires careful sampling of the transmitted bitstream during training to achieve scalable bitrates within a single model. This sampling strategy can lead to performance trade-offs across different bitrates and may cause instability during training. Therefore, future research in speech foundational models could explore leveraging alternative recurrent structures Toderici et al. (2017); Johnston et al. (2018); Diao et al. (2020) to improve coding scalability and address these challenges.
Appendix A Pre-training Paradigm
The proposed pre-training paradigm for optimizing vector quantization layers is detailed in Algorithm 3. During the pre-training phase, all vector quantization layers are bypassed, effectively reducing the codec to a standard autoencoder trained solely on reconstruction losses (Lines 2-4). Once the encoder and decoder reach a certain level of convergence, the VQ layers are reactivated, and joint optimization resumes. In the pre-training phase, we set the $\operatorname*{arg\,min}$ nearest-neighbor selection as an identity function, making $\bm{z}_{q}$ equal to the input vector $\bm{z}_{e}$.
Algorithm 3 Pre-training Paradigm
1:repeat
2:     $\hat{\mathcal{X}}=G_{\psi}(F_{\phi}(\mathcal{X}))$
3:     $\mathcal{L}=\mathcal{L}_{recon}(\mathcal{X},\hat{\mathcal{X}})$
4:     take gradient descent step on $\nabla_{\phi}\mathcal{L},\nabla_{\psi}\mathcal{L}$
5:until converged
6:activate VQs and continue learning as usual
Appendix B Experiment Details
B.1 DAC Reproduction Setups
Our customized reproduction of DAC models closely follows the official development scripts. The original DAC model, designed for 16kHz audio signals, employs 12 VQ layers in its residual VQ module, supporting bitrates ranging from 0.5 kbps to 6.0 kbps. To ensure a fair comparison with ESC at similar bitrate levels, we extended the number of VQ layers in the RVQ module to 18, resulting in DAC-Base. For DAC-Tiny, we reduced the encoder dimension from 64 to 32 and the decoder dimension from 1536 to 288, while keeping other parameters unchanged. All DAC baselines were trained for 0.4 million iterations with a batch size of 16 on our multilingual speech dataset. Additional configuration details can be found in the official release [^5] The official configuration for 16kHz DAC model is available at https://github.com/descriptinc/descript-audio-codec/blob/main/conf/final/16khz.yml.
B.2 ESC Architecture Configurations
Overall, all three ESC variants are trained in distributed setups for 0.4 million iterations across 4 NVIDIA RTX 4090 GPUs with a total batch size of 36. These experiments took approximately 100 GPU hours.
B.2.1 Model Parameters
The parameter configurations for ESC-Base are provided in Table 3. For STFT transformation, we use a 20 ms window length and a 5 ms hop length, implemented with torchaudio. The number of FFT points is set to 382, resulting in a frequency dimension of 192. In the Swin Transformer, the layer depth represents the number of Swin Transformer Blocks (STBs) cascaded at each encoder and decoder layer. We use GELU activation functions and LayerNorm for normalization. In the down-sampling/up-sampling module, we use a scaling factor of $v=2$ to un-shuffle/shuffle along the frequency resolution only. Before the vector quantization layers, ESC processes two overlapping time frames together. To implement this, the flattened spectrum feature $\mathcal{Z}$ is reshaped from $\mathbb{R}^{W_{i}\times H_{i}C_{i}}$ to $\mathbb{R}^{W_{i}/2\times 2H_{i}C_{i}}$. Each frame is then split into sub-vectors, down-projected, and $l_{2}$ normalized before computing the distance matrix. The VQ layer at each bitstream of ESC-Base consumes $\log_{2}1024\times 3\times 150=4500$ bits per 3-second input speech (i.e., 1.5 kbps bitrate). For the scaled-up ESC-Large variant, we increase the STB layer depth from 2 to 4 while keeping the other configurations unchanged.
Modules Parameters Values
STFT Window/Hop Length [20ms, 5ms]
Number of FFT 382
Encoder/Decoder Patch Size [3, 2]
Layer Dims $C_{1},...,C_{6}$ [45, 72, 96, 144, 192, 384]
Attention Heads [3, 3, 6, 12, 24, 24]
Layer Depth 2
Scaling Factor $v$ 2
Vector Quantization Product VQ Size $l$ 3
Codevector Dimension $u$ 8
Codebook Size $K$ 1024
Table 3: Parameter configurations of model variant ESC-Base, which comprises 6 encoder/decoder layers.
B.2.2 Adversarial Training Setup
In the ESC-Base (adversarial) variant, the GAN discriminator is identical to the one used in DAC, consisting of a multi-period discriminator (MPD), multi-band discriminator (MBD), and multi-scale STFT discriminator (MSD), totaling over 42 million parameters. The adversarial loss formulation follows the official DAC-Base (adversarial) configuration. Additionally, we maintain the pre-training paradigm for 0.75 million iterations in this variant, with the discriminator intervening in training only after the pre-training stage finishes.
B.3 Details on Ablation Experiments
All ablation models operate on the complex STFT spectrum, as in ESC (SwinT + CS-RVQ), using the same STFT configurations specified in Table 3. These models were trained for 0.25 million iterations, with 0.025 million iterations allocated for pre-training. The Swin Transformer configurations mirror those used in ESC-Base. Similarly, the vector quantization setup in the CS-RVQ models follows that of ESC-Base. In total, the ablation experiments required approximately 80 hours on 4 RTX 4090 GPUs.
B.3.1 Convolution Blocks
For models with CNN backbones, the convolutional channel dimensions were set to match the hidden sizes of the STB-based models. Each CNN block consists of one residual unit and one downsampling/upsampling 2D convolutional layer with a stride of 2 along the frequency resolution only. The residual unit consists of two 2D convolutional layers, each followed by BatchNorm and Parametric ReLU activation.
B.3.2 Residual Vector Quantization Setups
For models using RVQs, we adapted the basic RVQ framework commonly used in time-domain codecs. To process frequency-domain spectrum features at the latent bottleneck, we combined RVQ with product vector quantization. Specifically, the flattened time frame vector is split into sub-group vectors, which are then recursively quantized, as in standard RVQs. We set the number of product VQs to 3 and the number of residual VQs to 6, ensuring the bitrate levels match those of ESC-Base (1.5 kbps per bitstream, 6 in total).